OpenStudy (anonymous):

I need to use substitution method to find all solutions of the systems of equations for x^2 + y^2=25 and y=2x. I need help. The equations are in brackets.

OpenStudy (anonymous):

Substitute y=2x into the first equation. Then simplify it and you have a quadratic in x which you can solve by factorisation (possibly), completing the square or the formula.

OpenStudy (anonymous):

How do you work it out bfsgd? Help me please.

OpenStudy (anonymous):

Anyone!!! Help me.

OpenStudy (anonymous):

\[x^2 + 4x^2 = 25 \text{ so } x^2 = 5\]

OpenStudy (anonymous):

Ok. And then bfsgd.

OpenStudy (anonymous):

Then we take square roots. There are two solutions for x namely \[x = \pm \sqrt{5} \] and y=2x so \[y=\pm 2\sqrt{5}\]